If the matrix R =[-5 2]

[M -4]

is singular, determine the

value of m .


1
Expert's answer
2021-12-13T16:16:49-0500

A square matrix is singular if and only if its determinant is zero.


"R=\begin{bmatrix}
-5&2\\m&-4
\end{bmatrix}"


"\begin{vmatrix}
-5&2\\m&-4
\end{vmatrix}=(-5)\cdot (-4)-2m=20-2m=0"


Answer: "m=10" .